OpenStudy (anonymous):

obtain the differential equation of the family of curve and sketch several representative members of the family: Parabolas with axis parallel to x-axis

OpenStudy (anonymous):

my teacher already give the answer.. and this is the answer

OpenStudy (perl):

your teacher got 2y'* y'' - 3 y ''' = 0

OpenStudy (perl):

can you take a screenshot of that whole sheet

OpenStudy (anonymous):

none

OpenStudy (anonymous):

left side no.2

OpenStudy (perl):

since it says the parabola is parallel to the x axis (sorry my earlier work is incorrect)

OpenStudy (anonymous):

it's okay.. no need to apologize

OpenStudy (perl):

actually to be more general, since it says just parallel to the x axis (y-k)^2 = 4a (x-b)

OpenStudy (perl):

now we have three arbitrary constants, which we need to eliminate

OpenStudy (anonymous):

okay, the n next..

OpenStudy (perl):

(y-k)^2 = 4a (x-b) 2(y-k) * y ' = 4a * 1 2(y' ) * y' + 2(y-k) y' ' = 0

OpenStudy (anonymous):

y^2 - 2ky + k^2 = 4ax-4ab where k^2 and -4ab will become zero..

OpenStudy (anonymous):

am i right?

OpenStudy (perl):

yes,

OpenStudy (perl):

(y-k)^2 = 4a (x-b) y^2 - 2ky + k^2 = 4ax-4ab 2y y' - 2k y ' + 0 = 4a 2y' y' + 2y y' ' - 2k y ' = 0

OpenStudy (perl):

we still have that k, so we have to take derivative again

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

isn't it -ky''??

OpenStudy (anonymous):

-2ky''?

OpenStudy (perl):

yes

OpenStudy (perl):

(y-k)^2 = 4a (x-b) y^2 - 2ky + k^2 = 4ax-4ab 2y y' - 2k y ' + 0 = 4a 2y' y' + 2y y' ' - 2k y '' = 0 divide through by 2 , simplify (y ' )^2 + y * y ' ' - k y ' ' = 0

OpenStudy (anonymous):

transpose -ky'' to the right side then divide by y''??

OpenStudy (perl):

thats a great idea

OpenStudy (anonymous):

what's next??

OpenStudy (anonymous):

where??

OpenStudy (perl):

$$\Large{ (y ' )^2 + y \cdot y ' ' = k y ' ' \\ (y ' )^2 + y \cdot y ' ' = k y ' '\\ \frac{(y ' )^2}{y''} + \frac{y \cdot y ' '}{y''} = k \\ \frac{(y ' )^2}{y''} + y = k \\~\\ \frac{y ''~ 2(y ' )\color{red}{y ''} - (y')^2 y'''}{(y'')^2} + y' = 0 \\ }$$

OpenStudy (anonymous):

awhh, okay..

OpenStudy (perl):

then we can find a common denominator

OpenStudy (perl):

$$\Large{ (y ' )^2 + y \cdot y ' ' = k y ' ' \\ (y ' )^2 + y \cdot y ' ' = k y ' '\\ \frac{(y ' )^2}{y''} + \frac{y \cdot y ' '}{y''} = k \\ \frac{(y ' )^2}{y''} + y = k \\~\\ \frac{y ''~ 2(y ' )\color{red}{y ''} - (y')^2 y'''}{(y'')^2} + y' = 0 \\ \frac{y ''~ 2(y ' )\color{red}{y ''} - (y')^2 y'''}{(y'')^2} + y' \cdot \frac{(y'')^2}{(y'')^2}= 0 \\ }$$

OpenStudy (perl):

multiply both sides by ( y ' ' ) ^2

OpenStudy (anonymous):

in order to cancel the y''??

OpenStudy (anonymous):

in denominator??

OpenStudy (perl):

$$\Large{ (y ' )^2 + y \cdot y ' ' = k y ' ' \\ (y ' )^2 + y \cdot y ' ' = k y ' '\\ \frac{(y ' )^2}{y''} + \frac{y \cdot y ' '}{y''} = k \\ \frac{(y ' )^2}{y''} + y = k \\~\\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y'''}{(y'')^2} + y' = 0 \\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2}{(y'')^2 } = 0 \\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ 2y' (y ' ' )^2 - (y')^2 y ''' +y' \cdot (y'')^2=0\\~\\ 2y' (y ' ' )^2 +y' \cdot (y'')^2- (y')^2 y ''' = 0 \\~\\ 3y' (y ' ' )^2 - (y')^2 y ''' = 0 \\~\\ }$$

OpenStudy (perl):

can you read what your teacher wrote

OpenStudy (anonymous):

yes, y'(y''')-3(y'')^2=0

OpenStudy (perl):

we almost got that answer

OpenStudy (perl):

$$ \Large{ (y ' )^2 + y \cdot y ' ' = k y ' ' \\ (y ' )^2 + y \cdot y ' ' = k y ' '\\ \frac{(y ' )^2}{y''} + \frac{y \cdot y ' '}{y''} = k \\ \frac{(y ' )^2}{y''} + y = k \\~\\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y'''}{(y'')^2} + y' = 0 \\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2}{(y'')^2 } = 0 \\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ 2y' (y ' ' )^2 - (y')^2 y ''' +y' \cdot (y'')^2=0\\~\\ 3y' (y ' ' )^2 - (y')^2 y '''=0\\~\\ } $$

OpenStudy (perl):

oh divide by y '

OpenStudy (anonymous):

almost got it.. but i still wonder why this is the answer, i already solve this but still we both have the same answer..

OpenStudy (anonymous):

gotcha.. thank you so much..

OpenStudy (perl):

$$ \Large{ (y ' )^2 + y \cdot y ' ' = k y ' ' \\ (y ' )^2 + y \cdot y ' ' = k y ' '\\ \frac{(y ' )^2}{y''} + \frac{y \cdot y ' '}{y''} = k \\ \frac{(y ' )^2}{y''} + y = k \\~\\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y'''}{(y'')^2} + y' = 0 \\ \frac{y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2}{(y'')^2 } = 0 \\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ y ''~ 2(y ' ){y ''} - (y')^2 y''' + y' \cdot (y'')^2=0\\~\\ 2y' (y ' ' )^2 - (y')^2 y ''' +y' \cdot (y'')^2=0\\~\\ 3y' (y ' ' )^2 - (y')^2 y '''=0\\~\\ \frac{3y' (y ' ' )^2 - (y')^2 y '''}{y'}=\frac{0}{y'}\\~\\ 3(y ' ' )^2 - y' \cdot y '''=0 \\ ~\\\rm divide ~ by~ sides~ by -1\\~\\ y' \cdot y '''-3(y ' ' )^2=0 } $$

OpenStudy (anonymous):

me too, i wonder why...

OpenStudy (anonymous):

thank youu so much.,

OpenStudy (perl):

and to graph the family of curves you know how to do that .